# ViSA-Flow: Accelerating Robot Skill Learning via Large-Scale Video Semantic Action Flow

### v2

@@ 4.1 Simulation Experiments
Figure 3: Datasets used for pretraining, finetuning, and evaluation. A model is trained on the Something-Something-V2 dataset with text labels. Placeholders are used to extract underlying semantic action flow. The finetuning stage involves 34 manipulation tasks across three simulated environments (Env A, B, and C) in CALVIN benchmark [28]. The evaluation is on Environment D, where the robot complete 5 consecutive subtasks within one continuous sequence.
Pre-training Data. The ViSA-Flow model undergoes pre-training (Stage 1, Sec. 3.3) using the large-scale Something-Something-V2 (SthV2) dataset [29] as the source domain. SthV2 contains approximately 220,000 short videos depicting diverse human-object interactions (examples visualized in the upper row of Fig. 3). Each video is associated with a template-based textual description indicating the action performed (e.g., ‘Pushing [something] from left to right’) and includes placeholder labels identifying key objects within frames. The videos are processed to extract ViSA-Flow representations which are used for the pre-training as described in Secs. 3.2 and 3.3.
Fine-tuning Data. Following pre-training, ViSA-Flow is fine-tuned (Stage 2, Sec. 3.3) specifically for the CALVIN environment. To evaluate performance under data scarcity, we utilize only 10% (1,768 trajectories) of the available language-annotated robot demonstrations from CALVIN’s ABC dataset as our target domain dataset. Each trajectory consists of the language instruction and the sequence of robot states, observations, and actions.
@@
Method Fully-Annotated Data (Demo No.) Partially-Annotated Data Tasks Completed in A Row Avg. Len.
@@
Hulc [10] 100% (17870) 41.8% 16.5% 5.7% 1.9% 1.1% 0.67
MDT [17] 100% (17870) 61.7% 40.6% 23.8% 14.7% 8.7% 1.54
Spil [18] 100% (17870) 74.2% 46.3% 27.6% 14.7% 8.0% 1.71
Roboflamingo [19] 100% (17870) 82.4% 61.9% 46.6% 33.1% 23.5% 2.47
SuSIE [11] 100% (17870) 87.0% 69.0% 49.0% 38.0% 26.0% 2.69
CLOVER [8] 10% (1768) 44.3% 18.0% 5.0% 1.0% 0.0% 0.68
GR-1 [22] 10% (1768) 67.2% 37.1% 19.8% 10.8% 6.9% 1.41
SeeR [7] 10% (1768) 65.5% 38.8% 21.4% 11.7% 6.8% 1.44
GR-MG [6] 10% (1768) 81.8% 59.0% 39.0% 24.0% 16.2% 2.20
@@
Baselines. We compare ViSA-Flow against two groups of SOTA methods: (i) Low-Data Baselines: Strong contemporary methods trained under the identical 10% data condition as ViSA-Flow for direct comparison of data efficiency. This includes CLOVER [8], GR-1 [22], SeeR[7] and GR-MG [6]. (ii) Full-Data Baselines: Methods trained on 100% of CALVIN annotated robot data (17,870 trajectories), including Hulc [10], MDT [17], Spil [18], Roboflamingo [19] and SuSIE [11]. These represent the performance achievable with substantially more in-domain supervision.
@@
Metrics. Following the standard CALVIN evaluation protocol, we measure the success rate to complete 5 consecutive subtasks within a longer instruction sequence, evaluated over 1,000 independent sequences. We also report the average successful sequence length (Avg. Len.). These metrics assess single-task proficiency and the ability to maintain performance over long horizons.
@@
Results and Analysis. Table 1 presents the performance metrics for all methods. The results demonstrate that ViSA-Flow outperforms all baseline methods, achieving highest success rates across all consecutive task completion metrics despite using only 10% of the available annotated robot trajectories. Most impressively, ViSA-Flow maintains strong performance in sequential tasks, completing 5 consecutive tasks 31.4% of the time, almost twice the rate of the next best method trained with 10% data (GR-MG: 16.2%) and exceeding all methods trained on 100% data, including Susie (26.0%). The average sequence length of 2.96 further demonstrates the effectiveness of ViSA-Flow in handling long-horizon manipulation tasks. Performance degradation from single to sequential tasks (89.0% → 31.4%) is notably less severe for ViSA (64.7% reduction) compared to GR-MG (80.2% reduction) and Susie (70.1% reduction). This remarkable performance can probably be attributed to utilization of semantic action representations extracted from human demonstration videos. These results in simulation experiments validate our hypothesis that semantic action representations from human videos can significantly enhance robot skill learning, even when expert demonstrations are scarce and encounter different environments.
@@
Table 2: Ablation study evaluating the contribution of key components in ViSA-Flow.
@@
Ablation Study of ViSA-Flow Components. Table 2 summarizes the results when each component within the ViSA-Flow framework is individually removed from the full method. Removing the semantic entity grounding stage and tracking the motion of points across whole observation images significantly reduces performance across all consecutive-task metrics. Success rate on five-task sequences drops from 31.4% to just 9.6% with the average successful length falling from 2.96 to 1.64, which indicates the importance of accurately segmenting and identifying semantic entities to anchor tracking and flow conditioning. Omitting the robust temporal tracking stage decreases the average successful length over five-task sequences from 2.96 to 2.78, highlighting that consistent point correspondences are essential for preserving temporal dynamics across multi-step interactions. Excluding explicit manipulator grounding results in a modest drop in average sequence length, from 2.96 to 2.83, indicating that while segmentation and tracking are primary drivers of performance, manipulator cues still play a meaningful role in providing spatial context for action understanding. Overall, the full ViSA-Flow configuration—integrating segmentation, tracking, and manipulator grounding—achieves the best results across all metrics, confirming that each component contributes to capturing semantic action flow and enabling reliable long-horizon, cross-domain task execution.
4.2 Real World Experiments
@@
Experiment Setup. We evaluate our ViSA-Flow method in two real-world settings: two single-stage manipulation tasks and one long-horizon manipulation task. The demonstrations were collected by teleoperating a 7-DOF Franka Emika Panda arm using the Oculus-based application. We use two cameras (one eye-in-hand, one eye-to-hand) to provide RGB observations. The real-world experiment setup is shown in Fig. 4. For single-stage tasks, we collected 46 and 54 demonstrations for two tasks—MoveContainer and PickEggplant respectively. We train the ViSA-Flow policy for each single-stage task. For long-horizon tasks, we consider the same two subtasks, MoveContainer and PickEggplant, requiring the robot to complete the first task before sequentially solving the second. This setup ensures consistency with the testing scenario used in our simulation experiments. We evaluate each policy across 12 different initial positions.
Baselines. We compare our ViSA-Flow method with GR-MG [6] and the visuomotor Diffusion Policy (DP) [30], which leverages both RGB and proprioceptive inputs. To ensure fair comparison, all baseline models are trained on the same real-world demonstration datasets for the two single-stage tasks and the long-horizon task.
Figure 5: Real-world experimental results. Left: two single-stage tasks; Right: a two-stage long-horizon task.
@@
Figure 6: Qualitative results on the real world long-horizon task. We visualize the decoded ViSA-Flow prediction at ${\hat{z}}_{t+1}$ against the actual ViSA-Flow ${z}_{t+1}$ extracted from the next observation for four execution phases. Two rows correspond to the two subtasks that make up the long-horizon evaluation: (Top) Subtask 1 – MoveContainer. (Bottom) Subtask 2 – PickEggplant. Qualitatively, the model’s one-step predictions closely follow the true motion of the manipulator and task-relevant objects, even as the scene evolves across distinct interaction stages.
@@
5 Limitations and Future Work
While ViSA-Flow demonstrates strong performance in observational robot learning, it currently lacks explicit modeling of 3D geometry and contact dynamics, which may limit its generalization to tasks involving fine-grained physical interactions. The current framework also relies on pretrained VLM components that potentially restrict adaptability to novel domains. Future work includes enriching ViSA-Flow representations with contact physics and reducing reliance on pretrained components by jointly training ViSA-Flow with VLMs. Additionally, integrating ViSA-Flow’s priors with reinforcement learning algorithms and scaling pretraining to web-scale video corpora offer promising directions for advancing generalizable robot learning.
Appendix A Appendix
A.1 More Details about ViSA-Flow Representation
Desired Properties of ViSA-Flow.
We hypothesize that an effective ViSA-Flow representation $z_{t}=f(o_{t},l)$ should possess the following properties:
  1. 1.
    Encodes Interaction Structure: $z_{t}$ should explicitly encode the essential spatio-temporal relationships and relative motions between the primary manipulator (hand/gripper) and the key objects involved in the manipulation task described by $l$.
  2. 2.
    Invariance to Visual Nuisance Factors: The mapping $f$ should be robust to variations in $o_{t}$ that are irrelevant to the core interaction semantics, such as background clutter, lighting conditions, or specific object/manipulator textures. Formally, for a nuisance variation $\delta_{O}$, we desire $f(o_{t},l)\approx f(o_{t}+\delta_{O},l)$.
  3. 3.
    Cross-Domain Alignment of Task Structure: Despite morphological and appearance differences between source ($O^{S}$) and target ($O^{T}$) domains, many manipulation tasks share fundamental geometric and dynamic structures. ViSA-Flow should align these structures. If $o^{S}_{t}$ and $o^{T}_{t}$ depict the same semantic phase of a task (e.g., pre-grasp approach), their representations $z^{S}_{t}=f(o^{S}_{t},l)$ and $z^{T}_{t}=f(o^{T}_{t},l)$ should be proximate in $Z_{\text{ViSA-Flow}}$ under a suitable metric $d$, i.e., $d(z^{S}_{t},z^{T}_{t})\approx 0$.
Motivation for Transfer Learning via ViSA-Flow.
If ViSA-Flow exhibits these properties, particularly cross-domain alignment, then the underlying dynamics of a manipulation task, when modeled in the ViSA-Flow space $Z_{\text{ViSA-Flow}}$, should be more consistent across domains than in the raw observation spaces $O^{S},O^{T}$. Let $T_{task}:Z_{\text{ViSA-Flow}}\times A^{\prime}\rightarrow Z_{\text{ViSA-Flow}}$ represent these shared dynamics (where $A^{\prime}$ might be an abstract action space). A generative model $g_{\phi}(z_{t+1}|z_{\leq t},l)$ trained on sequences $\{z^{S}_{t}\}$ extracted from the large source dataset $D_{v}$ can learn a prior distribution capturing $T_{task}$. This learned prior, encoded in the parameters $\phi$, encapsulates structural knowledge about manipulation dynamics. When learning the target policy $\pi_{\theta}(a_{t}|z^{T}_{\leq t},s_{\leq t},l)$ using the limited target data $D_{\tau}$, initializing with or regularizing towards the pre-learned prior $g_{\phi}$ can significantly accelerate learning and improve data efficiency, as the model only needs to adapt the general dynamics to the specific target embodiment and refine the action mapping, rather than learning the dynamics entirely from the scarce target data.
A.2 Hyper-parameter Details
Table 3 lists the hyperparameters used in Section 3. The window sizes $h$, $k$, and $n$ set the length of recent ViSA-Flow representations, action chunk, and forward-prediction horizon, respectively. Loss weights $\lambda_{\text{fwd}}$ and $\lambda_{\text{prog}}$ balance action learning against auxiliary objectives; we down-weight the forward term during fine-tuning so the optimizer focuses on the action chunk prediction. Table 4 shows the hyper-parameters for training ViSA-Flow in different stages. Training (both pre-training and all fine-tuning) was performed on a single NVIDIA RTX 4090 GPU.
Table 3: Key hyper-parameters used for the design of ViSA-Flow architecture.
Hyper-parameter Pre-train Fine-tune (CALVIN) Fine-tune (Real world)
Predicted action length $k$ 5 10
Past observation length $h$ 10 10 10
Predicted ViSA-Flow length $n$ 3 3 3
ViSA-Flow loss ratio $\lambda_{\text{fwd}}$ 1.0 0.1 0.1
Progress loss ratio $\lambda_{\text{prog}}$ 1.0 1.0
Table 4: Hyper-parameters for pre-training and fine-tuning ViSA-Flow model.
Hyper-parameter Pre-train Fine-tune (CALVIN) Fine-tune (Real world)
Batch size 32 16 16
Base learning rate $3.6\times 10^{-4}$ $3.6\times 10^{-4}$ $3.6\times 10^{-4}$
Minimum LR scale $1\times 10^{-2}$ $1\times 10^{-2}$ $1\times 10^{-2}$
Weight decay 0.0 0.0 0.0
Optimizer Adam Adam Adam
Adam $\beta_{1}$ 0.9 0.9 0.9
Adam $\beta_{2}$ 0.999 0.999 0.999
Warm-up epochs 5 5 5
Training epochs 30 20 30 (single-stage), 50 (long-horizon)

### v3

@@ IV-A Simulation Experiments
Pre-training Data. The ViSA-Flow model undergoes pre-training (Stage 1, Sec. III-C) using the large-scale Something-Something-V2 (SthV2) dataset[goyal2017somethingsomethingvideodatabase] as the source domain. SthV2 contains approximately 220k short videos of human-object interactions with template text labels (examples visualized in the upper row of Fig. 3). The videos are processed to extract ViSA-Flow representations which are used for the pre-training as described in Secs. III-B and III-C.
Fine-tuning Data. Following pre-training, ViSA-Flow is fine-tuned (Stage 2, Sec. III-C) specifically for the CALVIN environment [mees2022calvin]. To evaluate performance under data scarcity, we utilize only 10% (1,768 trajectories) of the available language-annotated robot demonstrations from CALVIN’s ABC dataset as our target domain dataset. Each trajectory consists of the language instruction and the sequence of robot states, observations, and actions. Robot observations are processed into ViSA-Flow representations using the pipeline described in Sec. III-B.
@@
Method
Fully-Annotated
Data (Demo No.)
Partially-Annotated
Data
Tasks Completed in A Row Avg. Len.
@@
Hulc [mees2022matterslanguageconditionedrobotic] 100% (17870) 41.8% 16.5% 5.7% 1.9% 1.1% 0.67
MDT [reuss2024multimodaldiffusiontransformerlearning] 100% (17870) 61.7% 40.6% 23.8% 14.7% 8.7% 1.54
Spil [zhou2024languageconditionedimitationlearningbase] 100% (17870) 74.2% 46.3% 27.6% 14.7% 8.0% 1.71
Roboflamingo [li2024visionlanguagefoundationmodelseffective] 100% (17870) 82.4% 61.9% 46.6% 33.1% 23.5% 2.47
SuSIE [black2023susie] 100% (17870) 87.0% 69.0% 49.0% 38.0% 26.0% 2.69
ATM [wen2024anypointtrajectorymodelingpolicy] 10% (1768) 31.7% 5.1% 1.3% 0.0% 0.0% 0.43
CLOVER [bu2024closedloopvisuomotorcontrolgenerative] 10% (1768) 44.3% 18.0% 5.0% 1.0% 0.0% 0.68
GR-1 [wu2023unleashinglargescalevideogenerative] 10% (1768) 67.2% 37.1% 19.8% 10.8% 6.9% 1.41
SeeR [tian2024predictiveinversedynamicsmodels] 10% (1768) 65.5% 38.8% 21.4% 11.7% 6.8% 1.44
GR-MG [li2024grmg] 10% (1768) 81.8% 59.0% 39.0% 24.0% 16.2% 2.20
@@
Baselines.We compare ViSA-Flow against two groups of SOTA methods: (i) Low-Data Baselines trained on the same 10% split for a fair data-efficiency comparison—ATM [wen2024anypointtrajectorymodelingpolicy], a flow-based interface enabling any-point querying; CLOVER [bu2024closedloopvisuomotorcontrolgenerative], a generative closed-loop visuomotor controller; GR-1 [wu2023unleashinglargescalevideogenerative], a multimodal transformer pre-trained on human videos; SeeR [tian2024predictiveinversedynamicsmodels], a predictive inverse-dynamics approach; and GR-MG [li2024grmg], a closely related transformer that augments GR-1 with explicit goal conditioning. We chose these baselines because they span complementary representations (flow-based, generative modeling, inverse dynamics, language/goal-conditioned transformers) and include models both with and without human-video pretraining, allowing us to isolate the effect of cross-domain priors. (ii) Full-Data Baselines: Methods trained on 100% of CALVIN annotated robot data (17,870 trajectories), including Hulc[mees2022matterslanguageconditionedrobotic], MDT[reuss2024multimodaldiffusiontransformerlearning], Spil[zhou2024languageconditionedimitationlearningbase], Roboflamingo[li2024visionlanguagefoundationmodelseffective] and SuSIE[black2023susie]. These represent the performance achievable with substantially more in-domain supervision.
@@
Metrics. Following the standard CALVIN evaluation protocol[mees2022calvin], we measure the success rate to complete 5 consecutive subtasks, evaluated over 1,000 independent sequences. We also report the average successful sequence length (Avg. Len.). These metrics assess single-task proficiency and the ability to maintain performance over long horizons.
@@
Results and Analysis. Table I presents the performance metrics for all methods. The results demonstrate that ViSA-Flow outperforms all baseline methods, achieving highest success rates across all consecutive task completion metrics despite using only 10% of the available annotated robot trajectories. Most impressively, ViSA-Flow maintains strong performance in sequential tasks, completing 5 consecutive tasks 31.4% of the time, almost twice the rate of the next best method trained with 10% data (GR-MG: 16.2%) and exceeding all methods trained on 100% data, including Susie (26.0%). The average sequence length of 2.96 further demonstrates the effectiveness of ViSA-Flow in handling long-horizon manipulation tasks. Performance degradation from single to sequential tasks (89.0% → 31.4%) is notably less severe for ViSA (64.7% reduction) compared to GR-MG (80.2% reduction) and Susie (70.1% reduction). This remarkable performance can probably be attributed to utilization of semantic action representations extracted from human demonstration videos. These results in simulation experiments validate our hypothesis that semantic action representations from human videos can significantly enhance robot skill learning, even when expert demonstrations are scarce and encounter different environments.
Figure 5: Qualitative results on the real world long-horizon task. We visualize the decoded ViSA-Flow prediction at ${\hat{z}}_{t+1}$ against the actual ViSA-Flow ${z}_{t+1}$ extracted from the next observation for four execution phases. Two rows correspond to the two subtasks that make up the long-horizon evaluation: (Top) Subtask 1 – MoveContainer; (Bottom) Subtask 2 – PickEggplant. Qualitatively, the model’s one-step predictions closely follow the true motion of the manipulator and task-relevant objects, even as the scene evolves across distinct interaction stages.
Figure 6: Real‑world experimental results. Left: two single‑stage tasks; Right: a two‑stage long‑horizon task.
Ablation Study of ViSA-Flow Components. Table II summarizes the results when each component within the ViSA-Flow framework is individually removed from the full method. Critically, removing the human-video pre-training stage (w/o pre.) leads to a near collapse in performance, indicating that the dynamics prior distilled from large-scale human videos is essential for multi-step task success. Removing semantic entity grounding (w/o Seg., with $\alpha{=}0$) and tracking motion over whole images significantly reduces performance across all consecutive-task metrics: success on five-task sequences drops from 31.4% to 9.6%, and average successful length falls from 2.96 to 1.64, which indicates the importance of accurately segmenting and identifying semantic entities to anchor tracking and flow conditioning. Omitting temporal tracking (w/o Trace.) decreases average successful length from 2.96 to 2.78, highlighting that consistent point correspondences preserve temporal dynamics. Excluding manipulator grounding (w/o Hand) yields a modest drop (2.96 to 2.83) that validates segmentation and tracking are primary drivers while manipulator cues still aid spatial context. Overall, the full ViSA-Flow—integrating segmentation, tracking, and manipulator grounding—achieves the best results, and abtion results confirm that each component contributes to reliable long-horizon, cross-domain execution.
@@
TABLE II: Ablation study evaluating the contribution of key components in ViSA-Flow.
@@
TABLE III: Ablation study evaluating ViSA-Flow data scaling.
Data Tasks Completed in a Row Avg. Len.
1 2 3 4 5
[…]
Data Scaling. In addition, we evaluate ViSA-Flow under different amounts of robot demonstration data (5%, 10%, and 50%). As shown in Table III, performance consistently improves as the amount of data increases, highlighting ViSA-Flow’s scalability and data efficiency.
IV-B Real World Experiments
@@
Experiment Setup. We evaluate our ViSA-Flow method in two real-world settings: two single-stage manipulation tasks and one long-horizon manipulation task. The demonstrations were collected by teleoperating a 7-DOF Franka Emika Panda arm using the Oculus-based application. We use two cameras (one eye-in-hand, one eye-to-hand) to provide RGB observations. The real-world experiment setup is shown in Fig. 4. For single-stage tasks, we collected 46 and 54 demonstrations for two tasks—MoveContainer and PickEggplant respectively. We train the ViSA-Flow policy for each single-stage task. For long-horizon tasks, we consider the same two subtasks, MoveContainer and PickEggplant, requiring the robot to complete the first task before sequentially solving the second. This setup ensures consistency with the testing scenario used in our simulation experiments. We pre-train our ViSA-Flow model on an NVIDIA RTX 4090 for 30 epochs, followed by fine-tuning for 30 epochs on single-stage tasks and 50 epochs for long-horizon tasks, respectively. We evaluate each policy across 12 different initial positions.
Baselines. We compare our ViSA‑Flow method with GR‑MG [li2024grmg] and the visuomotor Diffusion Policy (DP) [chi2023diffusionpolicy], both of which leverage RGB and proprioceptive inputs. To ensure fair comparison, all baseline models are trained on the same real‑world demonstration datasets for the two single‑stage tasks and the long‑horizon task.
@@
V Limitations and Future Work
While ViSA-Flow demonstrates strong performance in observational robot learning, it currently lacks explicit modeling of 3D geometry and contact dynamics, which may limit its generalization to tasks requiring fine-grained physical interactions. The framework also relies on pretrained VLM components, potentially restricting adaptability to novel domains or unseen objects. Moreover, ViSA-Flow currently transfers only object–manipulator interactions from human manipulation videos to robots, ignoring more nuanced dexterous movements of human fingers. This simplification leads to some loss of manipulation knowledge, which may limit performance in tasks requiring fine or dexterous control.
Future work aims to address these limitations by enriching ViSA-Flow representations with contact physics and 3D reasoning, reducing reliance on pretrained models through joint or end-to-end training with VLMs, and integrating its learned priors with reinforcement learning to enhance policy learning. Additionally, we plan to investigate methods to capture and transfer finer-grained human manipulation skills, preserving dexterous finger-level knowledge for robotic use. Scaling pretraining to large-scale video corpora and further analyzing ViSA-Flow’s invariance properties and sample efficiency also represent promising directions for advancing robust and generalizable robot learning from observation.